### v4

# Half-Iterates and Delta Conjectures

Abstract
The vivid contrast between two competing algorithms for solving Abel’s equation $g(\theta(x))=g(x)+1$, given $\theta(x)$, is easily sketched.  EJ is faster and more efficient, but ML evaluates a limit characterizing the principal solution $g(x)$ directly.  EJ finds $g(x)+\delta$, where $\delta$ is possibly nonzero but independent of $x$.  If we were to know an exact expression for $\delta$, then the “intrinsicality” of ML would be subsumed by EJ.  Filling this gap in our knowledge is the aim of this paper.
00footnotetext: Copyright © 2025 by Steven R. Finch. All rights reserved.
The Mavecha-Laohakosol [1, 2, 3, 4] and Écalle-Jagy [5, 6, 7, 8] algorithms enable us to numerically solve functional equations
$\begin{array}[c]{ccc}g_{13}\left(x\left(1-x^{2}\right)\right)=g_{13}(x)+1,&&0<x<1;\end{array}$
$\begin{array}[c]{ccc}g_{14}\left(x(1-x+x^{2})\right)=g_{14}(x)+1,&&0<x<1;\end{array}$
$\begin{array}[c]{ccc}g_{15}\left(x(1-x)^{2}\right)=g_{15}(x)+1,&&0<x<1;\end{array}$
$\begin{array}[c]{ccc}g_{16}\left(x(1-3x+3x^{2})\right)=g_{16}(x)+1,&&0<x<1\end{array}$
in highly distinct ways.  It is useful to have both methods available.  For both, the base function $\theta(x)$ is assumed to be analytic with values $\theta(0)=0$ & $\theta^{\prime}(0)=1$; further, its Taylor series at the origin is
$\begin{array}[c]{ccccc}x+{\displaystyle\sum\limits_{m=1}^{\infty}}c_{m}x^{m\,\tau+1},&&c_{1}=\gamma<0,&&\tau\geq 1\text{ is an integer.}\end{array}$
In words, $\theta(x)$ possesses an attractive fixed point at $x=0$ with unit slope and, beyond this, series terms at uniform separation $\tau$.  Theory suggests that these conditions are sufficient for the ML algorithm to converge.  It seems that the same is true for the EJ algorithm, which is simpler than ML and outperforms ML in every regard.
Our approach is experimental, focusing mostly on cubics.  ML computes the principal solution $g(x)$ of Abel’s equation directly, whereas the EJ algorithm constructs $\lambda(x)$ that solves Julia’s equation
$\lambda(\theta(x))=\theta^{\prime}(x)\lambda(x)$
whose reciprocal approximates the derivative $g^{\prime}(x)$.  EJ has a drawback: it provides not $g(x)$ but instead $g(x)+\delta$, where $\delta$ is (usually) a nonzero constant.  Let $\theta_{13}(x)=x\left(1-x^{2}\right)$.  We wish to conjecture an exact expression for $\delta_{13}$, among other things.  Clearly $\tau=2$ and $\gamma=-1$.  EJ gives
$g_{13}(x)$ $=\frac{1}{2x^{2}}+\frac{3}{2}\ln(x)+\frac{5}{8}x^{2}+\frac{21}{32}x^{4}+\frac{35}{32}x^{6}+\frac{2717}{1280}x^{8}+\frac{13429}{3200}x^{10}+\frac{81239}{10752}x^{12}$
$+\frac{1271651}{125440}x^{14}+\frac{521379}{286720}x^{16}-\frac{18534919}{460800}x^{18}-\frac{1204285773}{11264000}x^{20}$
$+\frac{1901417927}{20815872}x^{22}+\frac{2501877529163}{1640038400}x^{24}+\cdots.$
To actually calculate $g_{13}(x)$ for some $0<x<1$, define $x_{0}=x$ and $x_{n}=x_{n-1}\left(1-x_{n-1}^{2}\right)$ for $n\geq 1$.  From the definition of $g_{13}(x)$, we have
$g_{13}(x)=\lim_{n\rightarrow\infty}\,\left[g_{13}(x_{n})-n\right].$
For example, to 100 decimal digits of accuracy,
$g_{13}\left(\frac{1}{2}\right)$ $=1.18935108880763516558234924956664105059667671443624\backslash$
        $\;\;\;\;\;\;\;\;71933977369986198458248324365808279005576819103340$...,
$g_{13}\left(\frac{1}{\sqrt{3}}\right)$ $=1.05481186044739525804034624653292094298561372154532\backslash$
        $\;\;\;\;\;\;\;\;98153825514557676637413068364378707969808390093840.$...
ML allows calculation of
$\tilde{g}_{13}(x)=-2\lim_{n\rightarrow\infty}n^{3/2}\left(\sqrt{2}\,x_{n}-\frac{1}{n^{1/2}}+\frac{3}{8}\frac{\ln(n)}{n^{3/2}}\right)$
to high precision and [9]
$\tilde{g}_{13}\left(\frac{1}{2}\right)$ $=1.70921147422759414764527334066027347665330181520643\backslash$
        $\;\;\;\;\;\;\;\;86339882470057398910413097076175322980529292243496$...,
$\tilde{g}_{13}\left(\frac{1}{\sqrt{3}}\right)$ $=1.57467224586735424010327033762655336904223882231552\backslash$
        $\;\;\;\;\;\;\;\;12559730614628877089577841074745751944760863233997.$...
The difference between the EJ-based & ML-based values is nonzero and conjectured to be
$\delta_{13}=g_{13}\left(1/2\right)-\tilde{g}_{13}\left(1/2\right)=g_{13}\left(1/\sqrt{3}\right)-\tilde{g}_{13}\left(1/\sqrt{3}\right)=-(3/4)\ln(2).$
This identity should hold for any choice of $x=x_{0}$.  A graph of $\tilde{g}_{13}\left(x\right)$ appears in Figure 1.  The half-iterate of $\theta_{13}(x)$ is defined by
$\begin{array}[c]{ccc}\theta_{13}^{[1/2]}(x)=g_{13}^{[-1]}\left(g_{13}(x)+\dfrac{1}{2}\right)&&\text{if }0<x<1\end{array}$
and is plotted in Figure 2.  For example,
$\theta_{13}^{[1/2]}\left(\frac{1}{2}\right)$ $=0.41820429029068859160564663263510996432355959664258\backslash$
        $\;\;\;\;\;\;\;\;55277260336535008451118441673117069901712981511710$...,
$\theta_{13}^{[1/2]}\left(\frac{1}{\sqrt{3}}\right)$ $=0.43400787521828089000086877815934526729705657496748\backslash$
        $\;\;\;\;\;\;\;\;10398836584355586827352894865816690477408133034496.$...
Generalizations include
$\delta(k)=-\frac{k+1}{2k}\ln(k)$
corresponding to $\theta(x)=x\left(1-x^{k}\right)$ for integer $k\geq 1$ and
$\delta(\rho)=\frac{1-\rho+\rho^{2}}{(1-\rho)^{2}}\ln\left(\frac{1}{1-\rho}\right)$
corresponding to $\theta(x)=x\left(1-x\right)(1+\rho\,x)$ for $0<\rho<1$.  For the former,
$\tilde{g}(x)=-k\lim_{n\rightarrow\infty}n^{(k+1)/k}\left(k^{1/k}\,x_{n}-\frac{1}{n^{1/k}}+\frac{k+1}{2k^{2}}\frac{\ln(n)}{n^{(k+1)/k}}\right)$
which resembles the case $k=2$ listed earlier.  For the latter,
$\tilde{g}(x)=-(1-\rho)\lim_{n\rightarrow\infty}n^{2}\left(\,x_{n}-\frac{1}{(1-\rho)n}+\frac{1-\rho+\rho^{2}}{(1-\rho)^{3}}\frac{\ln(n)}{n^{2}}\right)$
which does not resemble the case $\rho=1$ listed earlier.  Various other delta conjectures are listed in [8], building on work in [7].
We now turn to $\theta_{14}(x)=x(1-x+x^{2})$, for which $\tau=1$ and $\gamma=-1$.  EJ gives
$g_{14}(x)$ $=\frac{1}{x}-x-x^{2}-x^{3}-\frac{3}{4}x^{4}-\frac{1}{10}x^{5}+\frac{3}{4}x^{6}+\frac{13}{14}x^{7}-\frac{3}{4}x^{8}-\frac{167}{45}x^{9}$
$-\frac{57}{20}x^{10}+\frac{5081}{660}x^{11}+\frac{1837}{120}x^{12}-\frac{24235}{1092}x^{13}-\frac{14489}{168}x^{14}+\frac{269263}{3150}x^{15}+\cdots.$
To assess $g_{14}(x)$ for some $0<x<1$, define $x_{0}=x$ and $x_{n}=x_{n-1}\left(1-x_{n-1}+x_{n-1}^{2}\right)$ for $n\geq 1$.  ML allows calculation of
$\tilde{g}_{14}(x)=-\lim_{n\rightarrow\infty}n^{2}\left(\,x_{n}-\frac{1}{n}\right)$
to high precision and
$\tilde{g}_{14}\left(\frac{1}{2}\right)$ $=1.08497872758210320613109766431163045230197877946729\backslash$
        $\;\;\;\;\;\;\;\;58035039848467035694746977730608826554708937283892$...,
$\tilde{g}_{14}\left(\frac{1}{3}\right)$ $=2.51000182711894064597232155164267425915580827922384\backslash$
        $\;\;\;\;\;\;\;\;06758952435313433766715179785405812086576884568208.$...
We conjecture that $\delta_{14}=0$ and this holds for any choice of $x=x_{0}$.  A graph similar to that of $\tilde{g}_{14}\left(x\right)$ appears in [10], associated with $x(1-a\,x+x^{2})$ for $a=3/2$ (as opposed to $a=1$).  A generalization is $\delta(k)=0$ corresponding to both
$\begin{array}[c]{ccccc}x\,{\displaystyle\sum\limits_{j=0}^{k}}(-1)^{j}x^{j}&&\text{and}&&\dfrac{x}{1+x}\end{array}$
for integer $k\geq 1$; the limiting case $x/(1+x)$ is trivial [9].
We now turn to $\theta_{15}(x)=x(1-x)^{2}$, for which $\tau=1$ and $\gamma=-2$.  EJ gives
$g_{15}(x)$ $=\frac{1}{2x}+\frac{3}{4}\ln(x)+\frac{5}{8}x+\frac{21}{32}x^{2}+\frac{35}{32}x^{3}+\frac{2717}{1280}x^{4}+\frac{13429}{3200}x^{5}+\frac{81239}{10752}x^{6}+\frac{1271651}{125440}x^{7}$
$+\frac{521379}{286720}x^{8}-\frac{18534919}{460800}x^{9}-\frac{1204285773}{11264000}x^{10}+\frac{1901417927}{20815872}x^{11}+\frac{2501877529163}{1640038400}x^{12}$
$+\frac{779013297328117}{223865241600}x^{13}-\frac{10607531187899069}{803618816000}x^{14}-\frac{48418902026319109}{469647360000}x^{15}-\cdots.$
To assess $g_{15}(x)$ for some $0<x<1$, define $x_{0}=x$ and $x_{n}=x_{n-1}\left(1-x_{n-1}\right)^{2}$ for $n\geq 1$.  ML allows calculation of
$\tilde{g}_{15}(x)=-2\lim_{n\rightarrow\infty}n^{2}\left(\,x_{n}-\frac{1}{2n}+\frac{3}{8}\frac{\ln(n)}{n^{2}}\right)$
to high precision and [9]
$\tilde{g}_{15}\left(\frac{1}{2}\right)$ $=2.05147406138650942317849754883637246230944489958205\backslash$
        $\;\;\;\;\;\;\;\;11023132647875850901664278100401729946190648322822$...,
$\tilde{g}_{15}\left(\frac{1}{3}\right)$ $=1.57467224586735424010327033762655336904223882231552\backslash$
        $\;\;\;\;\;\;\;\;12559730614628877089577841074745751944760863233997.$...
We conjecture that $\delta_{15}=-(3/4)\ln(2)$ (identical to $\delta_{13}$) and this holds for any choice of $x=x_{0}$.  More generally,
$\delta(k)=-\frac{k+1}{2k}\ln(k)$
corresponding to $\theta(x)=x\left(1-x\right)^{k}$ for integer $k\geq 1$.  A graph of $\tilde{g}_{15}\left(x\right)$ appears in [10], associated with $x(1-a\,x+x^{2})$ for $a=2$.  The progression of graphs for $a=3/2$, $5/3$, $\sqrt{3}$, $9/5$ and $19/10$ is quite intriguing and deserves to be better known.
We finally turn to $\theta_{16}(x)=x(1-3x+3x^{2})$, for which $\tau=1$ and $\gamma=-3$.  EJ gives
$g_{16}(x)$ $=\frac{1}{3x}+\frac{2}{3}\ln(x)+\frac{2}{3}x+\frac{5}{6}x^{2}+\frac{14}{9}x^{3}+\frac{89}{30}x^{4}+\frac{563}{150}x^{5}-\frac{8749}{1260}x^{6}-\frac{22259}{294}x^{7}$
$-\frac{128419}{420}x^{8}-\frac{4596701}{9450}x^{9}+\frac{19562713}{8250}x^{10}+\frac{4523425099}{254100}x^{11}+\frac{495227461}{51480}x^{12}$
$-\frac{6781873063663}{13663650}x^{13}-\frac{171780306415997}{73573500}x^{14}+\frac{1764254963282093}{157657500}x^{15}+\cdots.$
To assess $g_{16}(x)$ for some $0<x<1$, define $x_{0}=x$ and $x_{n}=x_{n-1}\left(1-3x_{n-1}+3x_{n-1}^{2}\right)$ for $n\geq 1$.  ML allows calculation of
$\tilde{g}_{16}(x)=-3\lim_{n\rightarrow\infty}n^{2}\left(\,x_{n}-\frac{1}{3n}+\frac{2}{9}\frac{\ln(n)}{n^{2}}\right)$
to high precision and [9]
$\tilde{g}_{16}\left(\frac{1}{2}\right)$ $=1.11293077809135353800166857803390442442494189554367\backslash$
        $\;\;\;\;\;\;\;\;24343525392480517435909954503132620225124156973073$...,
$\tilde{g}_{16}\left(\frac{1}{3}\right)$ $=1.35456732398262548307271488105822274526124339313976\backslash$
        $\;\;\;\;\;\;\;\;49107656609639544183939435929936110800486434074608.$...
We conjecture that $\delta_{16}=-(2/3)\ln(3)$ and this holds for any choice of $x=x_{0}$.  Formal proofs would be good to see someday.  Updates will be reported in the Addendum.
Figure 1: Blue curve is $10\theta_{13}(x)$, scaled for visibility; green curve is ML-based $\tilde{g}_{13}(x)$. Lower left-hand corner is $(0,1)$. Distance between vertical notches is eight times that for horizontal.
Figure 2: Half-iterate of $x\left(1-x^{2}\right)$ in orange; $x\left(1-x^{2}\right)$ in blue; diagonal line in red.  Lower left-hand corner is $(0,1/10)$.
1 Clarification
A software programming error led to false conclusions in an earlier draft (v3).  The pseudocode in [7, 8] is entirely correct, but (inexplicably, over time) the expression ${\textstyle\sum\nolimits_{m=2}^{K-2}}v_{m}x^{\tau\,m+1}$ became altered to ${\textstyle\sum\nolimits_{m=3}^{K-1}}v_{m}x^{\tau\,m-1}$.  We mention this for completeness’ sake; every EJ-based result in that draft about $g_{14}$, $g_{15}$, $g_{16}$ is regrettably wrong.
2 Addendum
The quartic recurrence
$\begin{array}[c]{ccc}\xi_{n}=\dfrac{3}{4}+\dfrac{1}{4}\xi_{n-1}^{4},&&0\leq\xi_{0}<1\end{array}$
arises in the study of random Galton-Watson quaternary tree heights at criticality.  Letting $\xi_{n}=1-4x_{n}$, we obtain
$1-4x_{n}$ $=\frac{1}{4}\left[3+\left(1-4x_{n-1}\right)^{4}\right]$
$=\frac{1}{4}\left(4-16x_{n-1}+96x_{n-1}^{2}-256x_{n-1}^{3}+256x_{n-1}^{4}\right)$
$=1-4x_{n-1}+24x_{n-1}^{2}-64x_{n-1}^{3}+64x_{n-1}^{3}$
hence
$\begin{array}[c]{ccc}x_{n}=\theta_{17}(x_{n-1})=x_{n-1}\left(1-6x_{n-1}+16x_{n-1}^{2}-16x_{n-1}^{3}\right),&&0<x_{0}\leq 1/4.\end{array}$
ML allows calculation of
$\tilde{g}_{17}(x)=-6\lim_{n\rightarrow\infty}n^{2}\left(\,x_{n}-\frac{1}{6n}+\frac{5}{54}\frac{\ln(n)}{n^{2}}\right)$
to high precision and
$\tilde{g}_{17}\left(\frac{1}{4}\right)$ $=1.18801384280361604567470036016221695576085770053662\backslash$
        $\;\;\;\;\;\;\;\;68477703163061637332612940383107769158426699380380.$...
We conjecture that $\delta_{17}=-(5/9)\ln(6)$ and this holds for any choice of $x=x_{0}$.
3 Acknowledgements
I am grateful to Daniel Lichtblau at Wolfram Research for kindly answering my questions, e.g., about generalizing my original Mathematica code for ML to arbitrary $k$.  William Jagy [11] assisted me in more ways than he can imagine.  Dmitrii Kouznetsov [12] fleetingly mentioned both $g_{3}(\pi/2)$ and $\tilde{g}_{3}(\pi/2)$; see [7].  The creators of Mathematica earn my gratitude every day: this paper could not have otherwise been written.  An interactive computational notebook about ML is available [13] which might be useful to interested readers; an analog for EJ is forthcoming.

### v6

@@
 $\begin{array}[c]{ccc}g_{14}\left(x(1-x+x^{2})\right)=g_{14}(x)+1,&&0<x<1;\end{array}$
-$\begin{array}[c]{ccc}g_{15}\left(x(1-x)^{2}\right)=g_{15}(x)+1,&&0<x<1;\end{array}$
+$\begin{array}[c]{ccc}g_{15}\left(x(1-x)^{2}\right)=g_{15}(x)+1,&&0<x<2;\end{array}$
 $\begin{array}[c]{ccc}g_{16}\left(x(1-3x+3x^{2})\right)=g_{16}(x)+1,&&0<x<1\end{array}$
@@
 $+\frac{779013297328117}{223865241600}x^{13}-\frac{10607531187899069}{803618816000}x^{14}-\frac{48418902026319109}{469647360000}x^{15}-\cdots.$
-To assess $g_{15}(x)$ for some $0<x<1$, define $x_{0}=x$ and $x_{n}=x_{n-1}\left(1-x_{n-1}\right)^{2}$ for $n\geq 1$.  ML allows calculation of
+To assess $g_{15}(x)$ for some $0<x<2$, define $x_{0}=x$ and $x_{n}=x_{n-1}\left(1-x_{n-1}\right)^{2}$ for $n\geq 1$.  ML allows calculation of
 $\tilde{g}_{15}(x)=-2\lim_{n\rightarrow\infty}n^{2}\left(\,x_{n}-\frac{1}{2n}+\frac{3}{8}\frac{\ln(n)}{n^{2}}\right)$
@@
         $\;\;\;\;\;\;\;\;49107656609639544183939435929936110800486434074608.$...
-We conjecture that $\delta_{16}=-(2/3)\ln(3)$ and this holds for any choice of $x=x_{0}$.  Formal proofs would be good to see someday.  Updates will be reported in the Addendum.
+We conjecture that $\delta_{16}=-(2/3)\ln(3)$ and this holds for any choice of $x=x_{0}$.  Formal proofs would be good to see someday.  Updates will be reported in the Addenda.
 Figure 1: Blue curve is $10\theta_{13}(x)$, scaled for visibility; green curve is ML-based $\tilde{g}_{13}(x)$.
@@
 We mention this for completeness’ sake; every EJ-based result in that draft about $g_{14}$, $g_{15}$, $g_{16}$ is regrettably wrong.
-2 Addendum
+2 Addendum I
 The quartic recurrence
@@
 hence
-$\begin{array}[c]{ccc}x_{n}=\theta_{17}(x_{n-1})=x_{n-1}\left(1-6x_{n-1}+16x_{n-1}^{2}-16x_{n-1}^{3}\right),&&0<x_{0}\leq 1/4.\end{array}$
+$\begin{array}[c]{ccc}x_{n}=\theta_{17}(x_{n-1})=x_{n-1}\left(1-6x_{n-1}+16x_{n-1}^{2}-16x_{n-1}^{3}\right),&&0<x_{0}<1/2.\end{array}$
 ML allows calculation of
@@
 We conjecture that $\delta_{17}=-(5/9)\ln(6)$ and this holds for any choice of $x=x_{0}$.
+3 Addendum II
+The cubics in this paper are all of type $x\left(1-\sigma\,x-\rho\,x^{2}\right)$, where $\infty<\rho<\infty$ and $\sigma\geq 0$.  We conjecture, assuming $\sigma>0$, that
+$\delta=\left(1+\frac{\rho}{\sigma^{2}}\right)\ln\left(\frac{1}{\sigma}\right).$
+This extends an earlier formula connected with $0<\rho<1$ and $\sigma=1-\rho$.
+A more far-reaching conjecture starts with observations that both
+$\begin{array}[c]{ccccc}x(1-x)^{3}=x\left(1-3x+3x^{2}-x^{3}\right)&&\text{and}&&\theta_{16}(x)=x\left(1-3x+3x^{2}\right)\end{array}$
+share $\delta=-(2/3)\ln(3)$; both
+$\begin{array}[c]{ccccc}\theta_{17}(x)=x\left(1-6x+16x^{2}-16x^{3}\right)&&\text{and}&&x\left(1-6x+16x^{2}\right)\end{array}$
+share $\delta=-(5/9)\ln(6)$; both [7]
+$\begin{array}[c]{ccccc}\theta_{5}(x)=\ln(1+x)&&\text{and}&&x\left(1-\dfrac{1}{2}x+\dfrac{1}{3}x^{2}\right)\end{array}$
+share $\delta=-(1/3)\ln(2)$; both [8]
+$\begin{array}[c]{ccccc}\theta_{6}(x)=x\exp(-x)&&\text{and}&&x\left(1-x+\dfrac{1}{2}x^{2}\right)\end{array}$
+share $\delta=0$; both [8]
+$\begin{array}[c]{ccccc}\theta_{8}(x)=\dfrac{x}{1+x^{2}}&&\text{and}&&x\left(1-x^{2}+x^{4}\right)\end{array}$
+share $\delta=-(1/4)\ln(2)$; both [7]
+$\begin{array}[c]{ccccc}\theta_{3}(x)=\sin(x)&&\text{and}&&x\left(1-\dfrac{1}{6}x^{2}+\dfrac{1}{120}x^{4}\right)\end{array}$
+share $\delta=(3/5)\ln(3)$.  In essence, $\delta$ seems to be completely determined by the parameter $\tau$ and two coefficients $c_{1}$$c_{2}$.
+4 Addendum III
+Given a polynomial of type $x\left(1-\sigma\,x^{\tau}-\rho\,x^{2\tau}\right)$ where $\sigma>0$, we believe that
+$\delta=\left(\frac{\tau+1}{2\tau}+\frac{\rho}{\tau\,\sigma^{2}}\right)\ln\left(\frac{1}{\tau\,\sigma}\right).$
+This extends to transcendental functions, e.g., the Fresnel cosine [4] has $\delta=\frac{55}{108}\ln\left(\frac{10}{\pi^{2}}\right)$.
 5 Acknowledgements